OpenStudy (anonymous):

The legs of a right triangle are 9 cm and 40 cm long. What is the length of the hypotenuse?

OpenStudy (anonymous):

hey miss selena listen up... if you get your answer then plz give away a medal

OpenStudy (anonymous):

\[c=\sqrt(a^2+b^2)\] c=sqrt(1681) 41

OpenStudy (anonymous):

@Hashir she did give away a medal.

OpenStudy (anonymous):

41 = Hypotenuse

OpenStudy (anonymous):

^^she never do .... i solved 4 questions and she never gives one

OpenStudy (anonymous):

Don't know... she gave me a medal the last question.

OpenStudy (anonymous):

9^2 + 40^2 = H^2 Simplify. 81 + 1600 = H^2 Simplify. 1681 = H^2 Square root both sides. Hypotenuse = 41

OpenStudy (anonymous):

HOW DO I DO THAT ??